Circular No. 6927

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1855–026 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, reports on behalf of the RXTE ASM team at GSFC and the Massachusetts Institute of Technology: “An analysis of the RXTE All Sky Monitor light curve of the 362-s pulsar XTE J1855– 026 (IAUC 6904) obtained between 1996 Jan. 6 and 1998 May 28 reveals a modulation at a period of ∼ 6.1 days. A sine-wave fit yields 6.074 ± 0.004 days, with an amplitude of 0.15 counts/s and time of maximum flux of JD 2450289.6 ± 0.2. The mean count rate during this period was 0.5 counts/s, which corresponds to ∼ 7 mCrab (2–12 keV). This combination of pulse and orbital parameters raises the possibility that this may be a supergiant, rather than a Be star system, as earlier suggested. If this is a Be star system, then a 6-day orbital period would be the shortest known for this class.”

XTE J2012+381 N. E. White, Goddard Space Flight Center; and Y. Ueda, T. Dotani,

and F. Nagase, Institute of Space and Aeronautical Science, report: “This x-ray transient (IAUC 6920) was observed with the ASCA x-ray observatory from May 29.30 to 30.19 UT. The source position was determined to be α = 20^{h}12^{m}39^{s}.1, δ = +38^{o}10′50′′ (equinox 2000.0), with an uncertainty of 0′.5. This position contains that of the variable radio source (IAUC 6924). The ASCA position also overlaps the 1′ RXTE PCA error circle (IAUC 6922). The x-ray intensity in the band 0.7–4 keV increased linearly by 15 percent across the observation, whereas the 4–10-keV flux remained constant. The average 2–10-keV-band flux was 150 mCrab. Applying a disk blackbody and power-law model to the ASCA GIS spectrum in the range 1–10 keV yields an acceptable fit with a temperature of the innermost disk of 0.76 ± 0.01 keV, a power-law photon index of 2.9 ± 0.1, and an absorption column density (1.29 ± 0.03) × 10^{22} cm^{−2}. This ultrasoft spectrum with a hard power-law tail is the well established x-ray signature of blackhole candidates. If the x-ray absorption is due to the interstellar medium, it corresponds to 7 magnitudes of extinction in the V band.”

SUPERNOVA 1998bn IN NGC 4462 Visual magnitude estimates: May 5.76 UT, 13.5 (B. Monard, Pretoria,

S. Africa); 14.70, 14.0 (Monard); 23.8, 15.5 (L. Kiss, Ágasvár, Hungary).

1998 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green